Can you please help me do this problem? I have no example in my book.
2(x-1)squared=-16

Answer by stanbon(75887)

You need to understand complex numbers.
Divide by "2" to get (x+1)^2=-8
Take the square root to get:
x+1 = sqrt(-8)
x+1 = +-2sqrt(2)i
x= -1+2(sqrt2)i or x=-1-2(sqrt2)i
Cheers,
Stan H.